OpenStudy (anonymous):

. Choose the missing step in the given solution to the inequality -x - 10 > 14 + 2x. (5 points) -x - 10 > 14 + 2x -3x -10 > 14 x < -8 -3x > 24 -3x < 4 -3x > -4 8 < x

OpenStudy (amorfide):

-3x -10 > 14 if you rearrange this by adding 10 to both sides what do you get

OpenStudy (anonymous):

24

OpenStudy (amorfide):

what about the left hand side

OpenStudy (anonymous):

-3x>24

OpenStudy (amorfide):

now look at your answers given

OpenStudy (anonymous):

thanks

OpenStudy (amorfide):

glad to help

OpenStudy (anonymous):

can you help me with another one? @amorfide

OpenStudy (amorfide):

okay

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) 7x + 2(x - 1)? (5 points) A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 x x 3

OpenStudy (anonymous):

i have a hard time with these

OpenStudy (amorfide):

this is not an inequality as there is no > or < sign

OpenStudy (triciaal):

"is not a way" this means which choice is false and as @amorfide said you are missing the symbol in the original

OpenStudy (anonymous):

there is sorry it didnt show lol i will rewrite it

OpenStudy (anonymous):

5(x + 2) greater than or equal to 7x + 2(x - 1)

OpenStudy (amorfide):

okay so work it out for x first so you know what you are representing

OpenStudy (amorfide):

multiply out of the brackets, and get an inequality similar to x>23 (this is just an example of a random inequality )

OpenStudy (anonymous):

i dont understand sorry

OpenStudy (amorfide):

5(x + 2) greater than or equal to 7x + 2(x - 1) 5x+10 greater than or equal to 7x+2x-2 \[5x-9x \ge -2 -10\] \[-4x \ge -12\] \[x \le 3\]

OpenStudy (amorfide):

on a number line you would represent it with a filled in circle above the three (since it can be equal to 3) and then we want less than 3 also, so we shade in the part of the number line that is less than 3